Grade 11Physical Chemistry

volume occupied by one molecule of water (density=1g/cm3 is

amit
9 Years agoGrade 11

3 Answers

somi teez
9 Years ago
hi.
if we consider at STP VOLUME OCCUPIED BY ONE MOLE OR 6.023*10^23 molecules is 22.4 l
so,volume occupied by one molecule is 22.4/(6.023*10^23) l
saurabh gangwar
8 Years ago
we know that density of water is 1gm/ml.
molar vmass of water 18 gm/mole
number of moles in 1 gm =1/18=0.55 moles
18 gm contains 6.022×1023 molecules of h2o
 
no of molecule in one gram water 6.022×1023/18=0.33×1023
 
1 h2o molecule =1/0.33×1023
2.99×10-23
anindita
8 Years ago
6.022*10^23 molecules=18 gm of water
Therefore,1 molecule=18/6.022*10^23 gm of.                                                                              Water
                                      =3.0*10^-23gm of water
We know, density=mass/volume
                 Volume=mass/density
                               =3.0*10^-23gm/1gm cm^-3
                               =3.0*10^-23 cm^3